Home > Reports > 2006

Report Archive

2017 - 2016
2015 - 2014 - 2013 - 2012
2011 - 2010 - 2009 - 2008
2007 - 2006 - 2005 - 2004

Night Reports 2006

December 2006
Hours spent Observing: 166.4, Weather: 114.0, Technical downtime: 56.6

Date Hours Used Comment
ObservingWeatherTechnical
31 December10.90.00.0Photometric.
30 December10.90.00.0Photometric.
29 December7.90.03.0Photometric. Oil chiller fault (2.5), tracked into altitude limit (0.5).
28 December0.010.90.0Cloud, high humidity.
27 December0.010.90.0Thick cloud. TCS OK. Test observing at start of night (but frames mostly blank).
26 December0.05.45.5TCS computer failure. Cloudy second half.
25 December0.00.010.9TCS computer failure
24 December0.00.010.9TCS computer failure
23 December0.00.010.9TCS computer failure
22 December10.30.00.6Photometric. Alt lost, autorecovered ok.
21 December5.50.05.4Photometric. RCS problems (3 hours), Oil Chiller Problems (2.4 hours)
20 December0.010.90.0High humidity
19 December0.010.90.0Thick cloud, high humidity
18 December0.010.90.0Thick Cloud
17 December2.98.00.0Non-photometric then Rain.
16 December10.90.00.0Non-photometric.
15 December10.30.00.6Non-photometric. Alt axis limit.
14 December10.90.00.0Photometric.
13 December10.60.00.3Photometric. RCS problem.
12 December0.02.88.1Borderline atmospheric moisture count. Oil chiller problem (now fixed).
11 December0.010.90.0Ice, high humidity.
10 December0.010.90.0Freezing Fog, Ice buildup.
9 December0.010.80.0Cloud, freezing fog, ice.
8 December0.110.70.0Cloud, humidity, rain.
7 December10.80.00.0Non-photometric.
6 December10.80.00.0Photometric. Poor Seeing.
5 December10.80.00.0Photometric.
4 December10.40.00.4Photometric. Tracked into altitude limit.
3 December10.80.00.0Non-photometric.
2 December10.80.00.0Non-photometric.
1 December10.80.00.0Photometric.

November 2006

Hours spent Observing: 116.8, Weather: 171.4, Technical downtime: 29.5

Date Hours Used Comment
ObservingWeatherTechnical
30 November10.80.00.0Photometric.
29 November10.80.00.0Photometric.
28 November2.38.00.5Only north side of enclosure open, manual observing. Closed most of night by poor seeing. Autoguider deployment failure.
27 November10.70.00.0Autoguider testing and manual GRB observing, only North side of enclosure open.
26 November0.00.010.7Enclosure encoder fault.
25 November0.00.010.7Enclosure encoder fault.
24 November0.010.70.0High wind, humidity, ice.
23 November10.30.00.4Non-photometric. Auxiliary mechanism node fault, autorecovered.
22 November10.40.00.3Non-photometric. Altitude node fault, autorecovered.
21 November0.010.70.0High winds.
20 November0.010.60.0Thick cloud.
19 November0.010.60.0Thick cloud, high humidity
18 November0.010.60.0Thick cloud, high humidity
17 November10.60.00.0Non-photometric
16 November8.02.60.0High humidity at start of night, non-photometric
15 November0.010.60.0High humidity and rain.
14 November1.19.50.0Thick cloud, high humidity
13 November10.10.00.4Non photometric. Cass node lost, autorecovered ok.
12 November0.010.50.0Thick high cloud.
11 November0.010.50.0Thick high cloud.
10 November10.40.00.1Photometric. Tracked into rotator limit.
9 November10.50.00.0Photometric.
8 November0.04.46.0Thick cloud first half. On-mount power tripped.
7 November0.010.40.0High humidity.
6 November0.010.40.0Thick cloud.
5 November0.010.40.0Thick cloud.
4 November10.00.00.4Non photometric, especially second half. Cass node unresponsive, autorecovered.
3 November0.010.30.0Thick cloud.
2 November0.010.30.0Bad weather.
1 November0.010.30.0Bad weather.

October 2006

Hours spent Observing: 122.8, Weather: 152.1, Technical downtime: 7.8

Date Hours Used Comment
ObservingWeatherTechnical
31 October0.010.30.0Rain, wind, general nastiness.
30 October0.010.30.0Cloud
29 October0.010.20.0Bad weather
28 October0.010.20.0Bad weather
27 October0.010.20.0Bad weather
26 October0.010.20.0Wind, high humidity.
25 October0.010.10.0Cloud, rain, high wind.
24 October0.010.10.0Cloud, rain, high wind.
23 October7.52.60.0High Humidity at start of night. Photometric later.
22 October10.10.00.0Non-photometric.
21 October9.00.01.1Mountainwide mains supply failures, Non-photometric.
20 October9.20.80.0Thick cloud at start of night, then Non-photometrc.
19 October4.65.40.0High humidity at start of night, then Non-photometric
18 October0.010.00.0High humidity, rain.
17 October0.49.60.0High humidity most of night.
16 October0.09.90.0High humidity all night
15 October2.17.80.0Humidity, high winds. Non-photometric.
14 October9.00.00.9Non-photometric. RCS problem.
13 October9.80.00.0Photometric.
12 October5.50.00.34h scheduled maintenance, 0.3h CSC node dropout, autorecovered. Non-photometric.
11 October8.60.00.01.1 hours of scheduled maintenance.
10 October---Scheduled maintenance.
9 October---Scheduled maintenance.
8 October9.70.00.0Non-photometric.
7 October8.70.01.0Cassegrain axis fault, autorecovered.
6 October5.83.90.0Cloud, high winds.
5 October0.09.60.0Cloud, high winds.
4 October0.09.60.0Cloud, high winds.
3 October5.10.04.5Non photometric. Autoguider - RCS problems.
2 October9.50.00.0Non photometric.
1 October8.21.30.0Non photometric. High humidity.

September 2006

Hours spent Observing: 184.1, Weather: 33.6, Technical downtime: 44.0

Date Hours Used Comment
ObservingWeatherTechnical
30 September9.50.00.0Photometric.
29 September9.50.00.0Non-photometric.
28 September9.40.00.0Non-photometric.
27 September0.00.09.4SCC disk replaced. System rebuild underway
26 September0.00.09.4SCC disk failure. Engineer in transit to La Palma.
25 September0.00.09.3SCC disk failure.
24 September9.30.00.0Non-photometric.
23 September7.32.00.0High humidity at start of night. Non-photometric.
22 September3.41.24.6Non-photometric. AGS failure (4.6)
21 September0.68.60.0Non-photometric
20 September7.61.60.0Non-photometric. Cloudy first half.
19 September9.20.00.0Non-photometric. Second half of night probably was photometric.
18 September9.20.00.0Non-photometric
17 September0.09.10.0Cloud.
16 September9.10.00.0Photometric.
15 September9.00.00.0Photometric.
14 September9.00.00.0Photometric.
13 September4.52.71.8High humidity. RCS/OSS problem. Non-photometric.
12 September0.60.08.3RCS problem. Non-photometric.
11 September8.90.00.0Non-photometric.
10 September8.90.00.0Non-photometric.
9 September8.40.00.4Unresponsive nodes. Non-photometric
8 September8.80.00.0Non-photometeric.
7 September0.08.80.0Thick cloud
6 September8.20.00.5Cass rotator tracked into limit. Non-photometric
5 September8.70.00.0Non-photometric. Part autoguider tests.
4 September8.70.00.0Non-Photometric. Moderate Calima
3 September0.08.60.0Cloud, Rain, Calima.
2 September8.30.00.3Non-Photometric. AZM unresponsive, autorecovered.
1 September8.60.00.0Non-Photometric.

August 2006

Hours spent Observing: 191.5, Weather: 43.7, Technical downtime: 14.8

Headlines

  • Autoguider now in routine use on LT
  • WCS transformation provided for CCD data
  • Cloud monitor to be installed on site
  • Call for proposals for 2007A (PATT)
  • SupIRCam temporarily unavailable due to servicing
  • Technical downtime at a record low

Autoguider
We are pleased to report that an autoguider is now in successful routine use with the LT. Guided (closed loop) exposures may be made up to a maximum of 30 minutes; the overhead for use of the autoguider is 60 seconds. For short exposures (e.g. 2 minutes or less) it is generally recommended that unguided exposures are used since the telescope open-loop tracking suffices to provide good image quality.

WCS transformation provided for CCD data
WCS transformations are now written to the FITS header of all CCD data enabling WCS mapping of individual pixels of the array.

Cloud monitor to be installed on site
A cloud monitor has been procured and is currently undergoing testing. It is hoped to install it on site in the next few months.

Call for proposals for 2007A (PATT)
PATT has issued a call for proposals for the LT for semester 2007A. The deadline for applications is 30 September 2006. For more details see the PATT 07A news item.

SupIRCam temporarily unavailable due to servicing
Recently we have been experiencing increasing problems with SupIRCam, and the decision has been made to ship the camera back to the UK for major overhaul. Accordingly SupIRCam will be unavailable for observations for the next several months. Every effort is being made to get SupIRCam back online as soon as feasible; we will keep you informed of progress.

Technical downtime at a record low
The average operational availability of the telescope since the last report is 90.1%, just surpassing our target of 90%. For August, the telescope operational availability equalled our best previous record of 93% of time and the telescope was on-line, observing completely unsupervised for over 191 hours making this our current record month.

Date Hours Used Comment
ObservingWeatherTechnical
31 August8.50.00.0Non-photometric.
30 August0.08.50.0Dense Calima (dust) & Cloud.
29 August0.08.50.0Dense Calima (dust) & Cloud.
28 August8.50.00.0Non-photometric. Calima.
27 August8.40.00.0Possibly Photometric, some calima
26 August8.40.00.0Photometric
25 August8.40.00.0Photometric
24 August8.00.00.3Azimuth node unresponsive, autorecovered. CCD 10 degrees warmer than nominal.
23 August0.50.07.8Power supply interruptions tripped circuit breakers
22 August8.30.00.0Photometric. Part autoguider/RCS test
21 August8.20.00.0Photometric. Part autoguider/RCS tests
20 August8.20.00.0Photometric.
19 August8.20.00.0Photometric.
18 August3.12.92.1Non-photometric. High humidity. RCS autoguider problem, azimuth node unresponsive, autorecovered.
17 August0.08.10.0Cloud. GRB alert test.
16 August0.08.10.0Cloud
15 August4.00.04.0RCS autoguider problem.
14 August8.00.00.0Non-photometric. Patchy cloud. RCS problem.
13 August7.40.00.6Non-photometric. TCC(0.1) and AMN(0.5) unresponsive. Auto recovered
12 August7.90.00.0Photometric
11 August7.90.00.0Photometric
10 August7.90.00.0Photometric
9 August7.80.00.0Photometric
8 August7.80.00.0Photometric
7 August7.80.00.0Photometric
6 August7.70.00.0Photometric
5 August7.70.00.0Photometric. Windy
4 August7.70.00.0Photometric
3 August0.07.60.0Calima
2 August7.60.00.0Non-photometric. Mild calima
1 August7.60.00.0Photometric.
July 2006

Hours spent Observing: 187.8, Weather: 8.4, Technical downtime: 27.1

Date Hours Used Comment
ObservingWeatherTechnical
31 July7.50.00.1Photometric. TCC unresponsive, autorecovered.
30 July7.50.00.0Photometric
29 July3.60.03.9Elc node unresponsive, autorecovered. Fold mirror rotation problem. Many failed exposures.
28 July0.00.07.5Power failure to instrument supply
27 July6.40.01.0Robotic software corrupted. Photometric
26 July7.40.00.0Photometric
25 July7.40.00.0Non-photometric - Thinner Calima
24 July0.07.30.0Calima
23 July1.00.06.3Main Oil Pump Tripped, Calima
22 July6.21.10.0Non-photometric. Humidity spikes. Calima
21 July7.30.00.0Photometric.
20 July7.30.00.0Photometric.
19 July7.20.00.0Photometric.
18 July7.20.00.0Photometric.
17 July7.20.00.0Non-photometric.
16 July7.20.00.0Non-photometric.
15 July7.10.00.0Photometric.
14 July7.10.00.0Photometric.
13 July7.10.00.0Photometric.
12 July0.00.07.1Mountain wide electricity supply problems
11 July7.10.00.0Photometric.
10 July7.10.00.0Photometric.
9 July6.70.00.3Non-photometric. Unresponsive node autorecovered.
8 July7.00.00.0Non-photometric. Substantial cloud in 2nd half of night.
7 July7.00.00.0Non-photometric. Just a little cirrus around at dawn.
6 July7.00.00.0Photometric.
5 July6.70.00.3Photometric. Unresponsive node autorecovered.
4 July6.70.00.3Photometric. Az Ven failure, autorecovered.
3 July6.90.00.0Photometric.
2 July6.90.00.0Photometric.
1 July6.60.00.3Photometric. Az Ven failure, autorecovered.

June 2006

Hours spent Observing: 122.6, Weather: 3.6, Technical downtime: 28.6

Engineering upgrade work
The telescope was taken offline for a period of approximately 8 days for major engineering work in mid-June. The telescope is now back online. The outcome of the engineering work is as follows:

  • The azimuth tape encoder was replaced with a new tape. This has resulted in a very marked improvement in open-loop tracking for the telescope.
  • An autoguider and associated optics were installed on the telescope. Closed-loop tracking with this instrument has been achieved. Further work is in progress to integrate the instrument with the robotic control system of the telescope. This work is on schedule to ensure provision of a fully functional system for users at the start of the coming semester (2006B).
  • The fractured coolant pipes for SupIRCam were replaced by new pipes of a more flexible material. Unfortunately there is a, as yet unexplained, slow leak of coolant from the camera. This is being investigated as a matter of urgency. Users will be notified when this problem has been solved; in the meanwhile, SupIRCam will remain offline.
Time allocations for semester 2006B

All three allocation committees (PATT,CAT and (most recently) JMU) have met and assigned time allocations for proposals in semester 2006B. PIs for JMU proposals should receive feedback and time allocations shortly.

We are now receiving PHASE 2 forms for 2006B. Please ensure these are submitted as soon as possible to ensure no delay to the start of your programme at the beginning of the semester (1 August 2006).

There are only two weeks remaining for semester 2006A. If you have remaining observations to be done, be sure to submit the appropriate Phase 2 form as soon as possible though for the reasons given above, SupIRCam will not be available for the remainder of the semester.

Enhanced features for the Telescope system
  • Quick Look facility : All data taken with the Liverpool Telescope is now available online within approximately 15 minutes of the observation. Note carefully that the quick look data do not use the most up to date flat fields and have not undergone any quality control inspection. Full reductions using the most recent flat fields and which are passed by quality control inspection are typically available from the data archive and Recent Data resources the following day. For further details, see: quicklook.
  • Tracking for Near-Earth objects has improved : At users' request we recently changed the way we handle observations of rapidly moving targets to make it easier to observe Near-Earth objects. An excellent opportunity to test the new software came with the close pass of 2004XP14 last week which was enthusiastically covered by much of the media. On the night of 3rd July, the asteroid was moving at a non-siderial rate of 2 arcminutes per minute and the robotic scheduler was able to successfully make observations.
Telescope Reliability
In the period since the last newsletter (approximately 2 months) the telescope usage figures are: observing, 407.4 hours; time lost to weather, 113.8 hours; and time lost to technical faults (excluding scheduled maintainance period), 116.1 hours. The technical downtime was 22% of available time, and exceeded our target figure of 10%. Much of this downtime was due to errors associated with the azimuth tape encoder system, and this contributed to the decision to replace the tape. Now that a new tape has been successfully installed (see 'Engineering Upgrade Work' above), it is anticipated that technical downtime will be considerably reduced, and initial figures for July are indeed encouraging; so far only 3.5%.

National Schools' Observatory (NSO)
We are pleased to announce that thanks to sponsorship from LJMU, schools can now make observations with the Liverpool Telescope free through the NSO. We are now busy developing many new resources for schools to use, so please encourage every teacher you know to register at www.schoolsobservatory.org.uk/register !

Open Day on La Palma
A very successful open day was held at the La Palma Observatory on Saturday 8 July. The Liverpool Telescope had some 400 visitors who watched the telescope being controlled remotely from LJMU in the UK.

Date Hours Used Comment
ObservingWeatherTechnical
30 June5.90.01.0Photometric. Autorecovery problem following Az Ven failure.
29 June1.00.05.9Non-photometric. Network link to La Palma lost. Contact lost while in manual engineering mode.
28 June6.90.00.0Photometric. CCD camera offline.
27 June6.90.00.0Photometric.
26 June6.60.00.3Photometric. Azimuth ven failure; autorecovered. Alt returned to full range movement.
25 June6.60.00.3Photometric. Azimuth ven failure; autorecovered. Restricted Alt range.
24 June6.90.00.0Photometric. Restricted Alt range.
23 June2.60.04.3Photometric. Autorecovery problems following Ven failure. Restricted Alt range.
22 June6.90.00.0Photometric. Autoguider comissioning at start of night. Restricted Alt range.
21 June6.90.00.0Photometric. Restricted Alt range.
20 June3.40.03.5Photometric. Autorecovery problems following Ven failure. Restricted Alt range.
19 June6.90.00.0Photometric. Restricted Alt range.
18 June5.31.60.0Intervals of high humidity. Restricted Alt range.
17 June6.90.00.0Photometric. Restricted Alt range.
16 June6.70.00.2Photometric. SCC problem. Restricted Alt range. Az returned to full free movement.
15 June5.90.00.0Scheduled maintenance at start of night. Restricted Az,Alt ranges.
14 June3.00.00.0Non-photometric. Half night of scheduled maintenance. Restricted Az,Alt ranges.
13 June---Scheduled Maintenance
12 June---Scheduled Maintenance
11 June---Scheduled Maintenance
10 June---Scheduled Maintenance
9 June---Scheduled Maintenance
8 June---Scheduled Maintenance
7 June7.00.00.0Photometric, Restricted range in Alt and Az.
6 June0.02.05.0Humidity, Focus Problems
5 June7.00.00.0Restricted range in Alt and Az.
4 June0.00.07.0AMN node unresponsive. Focus problems after restart.
3 June6.60.00.5Photometric. Restricted range in Alt and Az. Tracked into Az limits (0.5)
2 June6.60.00.5Non-photometric. Restricted range in Alt and Az. Tracked into Az limits (0.5)
1 June7.00.00.1Non-photometric. Restricted range in Alt and Az. Tracked into Az limits (0.1)

May 2006

Hours spent Observing: 141.1, Weather: 23.7, Technical downtime: 66.1

Date Hours Used Comment
ObservingWeatherTechnical
31 May6.60.00.5Photometric. Restricted range in Alt and Az. Tracked into Az limits (0.5)
30 May6.80.00.3Photometric. Restricted range in Alt and Az. Azimuth node dropped off network, autorecovered in 0.3h
29 May7.10.00.0Photometric. Restricted range in Alt and Az.
28 May6.90.00.2Photometric. Restricted range in Alt and Az. Tracked into Az limits (0.2)
27 May7.10.00.0Photometric. Restricted range in Alt and Az.
26 May0.00.07.2Camera controller fault.
25 May6.70.00.5Photometric. Restricted range in Alt and Az. Tracked into limits (0.5)
24 May7.10.00.2Non-Photometric. Restricted range in Alt and Az. Tracked into limits (0.2)
23 May7.30.00.0Photometric. Restricted range in Alt and Az.
22 May0.00.07.3Scheduler problem.
21 May1.00.06.3Az encoder errors(1.1); Oil level fault (5.2).
20 May3.10.04.2Az encoder errors(3.7); Oil level fault (0.5).
19 May1.90.05.4Az encoder errors(1.1); RCS restart in engineering mode after rack reboot (4.3).
18 May0.00.07.4Oil level fault.
17 May5.10.02.3Non-photometric. Operating over restricted Alt range. AZM encoder errors.
16 May0.07.40.0Thick cloud.
15 May5.00.52.0High humidity at start of night. Non-Photometric. Operating over restricted Alt range. AZM encoder errors.
14 May6.10.01.4Photometric. Operating over restricted Alt range. AZM encoder errors.
13 May6.00.01.5Photometric. Operating over restricted Alt range. AZM encoder errors.
12 May5.30.02.2Non-Photometric. Operating over restricted Alt range. AZM encoder errors.
11 May6.10.01.5Photometric. Operating over restricted Alt range. AZM encoder errors.
10 May6.30.01.3Photometric. Operating over restricted Alt range. AZM encoder errors.
9 May2.80.04.8Non-photometric. Operating over restricted Alt range. AMN unresponsive and did not recover.
8 May6.80.00.8Non-photometric. Operating over restricted Alt range. AZM encoder errors
7 May6.70.01.0Non-photometric. Operating over restricted Alt range. AZM encoder errors
6 May5.60.02.1Photometric. Operating over restricted Alt range. AZM encoder errors
5 May6.40.01.3Photometric. Operating over restricted Alt range. AZM encoder errors
4 May7.00.00.8Photometric. Operating over restricted Alt range. AZM encoder errors
3 May4.30.03.6Photometric. Operating over restricted Alt range. AZM encoder errors
2 May0.07.90.0Humidity
1 May0.07.90.0Humidity/Rain

April 2006

Hours spent Observing: 143.7, Weather: 86.5, Technical downtime: 21.4

Date Hours Used Comment
ObservingWeatherTechnical
30 April7.00.00.9Photometric. Operating over restricted Alt range. AZM encoder errors (0.9)
29 April6.60.01.3Photometric. Operating over restricted Alt range. AZM encoder errors (1.3)
28 April7.00.01.0Photometric. Operating over restricted Alt range. AZM encoder errors (1.0)
27 April7.00.01.0Non-photometric. Operating over restricted Alt range. ALT flywheeling (0.6) AZM encoder error (0.4)
26 April8.00.00.0Non-photometric. Operating over restricted Alt range.
25 April0.00.08.1Enclosure failed to open.
24 April0.06.22.0High humidity for most of night. Enclosure failed to open.
23 April6.70.01.5Photometric. AZM lost homing, autorecovered (1.5). Operating over restricted Alt range.
22 April6.60.01.6Photometric. AZM lost homing, autorecovered (1.6). Operating over restricted Alt range.
21 April0.08.20.0Humidity
20 April7.00.21.0Non-photometric. AZM lost homing x 2 , autorecovered (0.6). Cas unresponsive, autorecovered (0.4). Operating over restricted Alt range.
19 April8.10.00.1Photometric. AZM lost homing, autorecovered. Operating over restricted Alt range.
18 April0.08.30.0Humidity
17 April6.51.80.0Thick cloud at start of night.
16 April0.08.40.0Thick Cloud
15 April7.70.00.7Non-photometric. AZM lost homing, autorecovered ok.
14 April8.50.00.0Non-photometric.
13 April0.08.60.0Humidity
12 April0.08.60.0Humidity
11 April6.62.00.0Cloud
10 April1.17.50.0Humidity until last hour of night
9 April0.08.60.0Wind & Rain.
8 April1.07.60.0Wind & Rain.
7 April8.60.00.0Photometric. Operating over restricted Alt range due to azimuth instability.
6 April7.10.01.3Photometric. Operating over restricted Alt range due to azimuth instability. AZM unresponsive, autorecoved ok.
5 April7.90.00.9SupIRCam only. Non-photometric. Operating over restricted Alt range due to azimuth instability.
4 April7.11.70.0High winds at start of night. Operating over restricted Alt range due to azimuth instability.
3 April0.08.80.0High winds.
2 April8.80.00.0Non-Photometric. Operating over restricted Alt range due to azimuth instability.
1 April8.80.00.0Photometric. Operating over restricted Alt range due to azimuth instability.

March 2006

Hours spent Observing: 175.7, Weather: 51.5, Technical downtime: 35.8

Reliability and Downtime
Telescope reliabilty has been good through February and March, with less than 10% downtime. Over the two months, three full nights have been lost to three different unrelated errors. All were quickly resolved once we had someone on site the following day. This is the expected nature of the telescope. All other faults have been corrected remotely during the day before start-up or automatically by the software during the night.

Engineering
We had a major engineering visit this month involving several staff and installation of a variety of new facilities and fixes.

  • The broken cryogenic system on SupIRCam has been restored and the instrument is currently operating. Its performance is still under daily review, but for the time being at least, it is in routine nightly operation. Users with SupIRCam time are encouraged to submit their phase 2 forms.
  • The new encoder system developed by TTL to deal with an alignment problem with the secondary mirror has been commissioned and tested. Pointing and open-loop tracking performance have both been significantly enhanced by this. The blind pointing accuracy is currently 6 arcsec RMS and we anticipate this will improve further in the coming months.
  • Following the unsuccessful work late last year, a second attempt was made at correcting the astigmatism in the primary mirror cell. The new arrangement of pneumatic actuators in the mirror cell appears to have worked very well. Performance will continue to be monitored over the coming months to ensure that this has genuinely been a long term solution.
  • A change to mounting procedures for the A&G box has led to a significant increase in the repeatability and evenness of motion in the cassegrain derotator. This long standing problem, though not appearing as any obvious fault in the science data, has been degrading pointing performance.
  • A change to the network wiring infrastructure has reduced the number of network communications errors (appearing as "node unresponsive" in the night logs), but not yet fully resolved this issue.

User Support and Time Allocation
A change has been made this month to the way the robotic control system handles exceptionally poor seeing. If the system's automated assessment of the seeing degrades past 3 arcsec, the RCS will revert to a background mode and will simply monitor photometric standard fields until it judges the conditions are sufficient to recommence science operations. This means that on rare, very poor nights you might see a large number of standards being taken.

The PATT 06B application round yielded an over-subscription rate of 1.7. The panel plans to meet in mid-May. The JMU TAC application deadline will be the end of April for semester 06B.

Among the new facilities aleady installed, tested and available for semester 06B are Ephemeris Groups and the Ringo instrument. See the phase 1 web pages for more details.

Autoguider
As described in our last report, we are working to an ambitious schedule to deliver the new autoguider for the telescope. The required mechanical work is now almost complete. There is a small amount of electrical work to do, but the main task now is software development, which is still on schedule for installation during the Summer as was originally planned.

Date Hours Used Comment
ObservingWeatherTechnical
31 March8.50.50.0Non-photometric. Operating over restricted Alt range due to azimuth instability.
30 March5.33.70.0Non-photometric. High winds. Operating over restricted Alt range due to azimuth instability.
29 March8.70.00.3Non-photometric. Operating over restricted Alt range due to azimuth instability. AMN unresponsive/reboot (0.3)
28 March0.00.09.0Instrument Control Computer failed
27 March9.10.00.0Non-photometric. Operating over restricted Alt range due to azimuth instability.
26 March7.02.10.0Thick cloud. Non-photometric. Operating over restricted Alt range due to azimuth instability.
25 March9.10.00.0Non-photometric. Operating over restricted Alt range due to azimuth instability.
24 March0.00.09.2Azimuth instability and oil distribution fault
23 March4.60.04.6Non-photometric. Problems with Azimuth axis instability
22 March9.20.00.0Non-photometric.
21 March9.00.20.0Non-photometric. 3 hours of scheduled engineering at start of night.
20 March9.30.00.0Photometric
19 March5.34.00.0High humidity at start and end of night
18 March8.21.10.0High humidity during second half of night
17 March0.09.40.0High winds and rain all night.
16 March0.09.50.0High winds and rain all night.
15 March5.63.50.4Non-photometric. High winds. AZM lost homing, autorecovered ok.
14 MarchScheduled engineering
13 MarchScheduled engineering
12 MarchScheduled engineering
11 March9.60.00.0Photometric
10 March5.20.04.4Photometric. Power distribution fault.
9 March8.60.01.0Photometric. CCD ontrol computer locked. Manually rebooted.
8 March9.10.00.5Non-Photometric. Calima, not cloud. AMN unresponsive, autorecovered (0.5)
7 March9.70.00.0Photometric.
6 March9.70.00.0Photometric. 5" seeing
5 March9.70.00.0Photometric
4 March0.03.76.0High humidity, Rain sensor problems
3 March5.84.00.0High humidity from 01:30
2 March9.40.00.4Non photometric, AMN unresponsive, autorecovered ok.
1 March0.09.80.0Ice

February 2006

Hours spent Observing: 128.8, Weather: 139.4, Technical downtime: 16.4

Date Hours Used Comment
ObservingWeatherTechnical
28 February0.09.80.0Snow and Ice, High humidity
27 February0.09.80.0High humidity, rain
26 February9.90.00.0Non-photometric
25 February0.09.90.0Humidity
24 February9.60.00.3Reboot for encoder errors.
23 February8.51.50.0Non-photometric. Cloud inceasing through night.
22 February10.00.00.0Photometric
21 February0.00.010.0Power supply to P36 Canbus node
20 February10.00.00.0Photometric.
19 February10.00.00.0M2 encoder commissioning. Non-photometric.
18 February3.53.53.0M2 encoder commissioning, pointing problems, high humidity
17 February10.10.00.0Photometric.
16 February8.00.02.1Photometric. RCS upgrade lost 2.1 hours
15 February0.010.10.0Ice
14 February0.010.10.0Freezing fog
13 February0.010.20.0High winds, thick cloud
12 February0.010.20.0High humidity all night
11 February0.010.20.0High humidity all night
10 February0.010.30.0High humidity all night
9 February0.010.30.0Very wet and stormy. Snow.
8 February0.010.30.0Very wet and stormy. Snow.
7 February0.010.40.0Very wet and stormy
6 February0.010.40.0Humidity
5 February10.40.00.0Non-photometric
4 February10.40.00.0Non-photometric. (Very very little cloud)
3 February10.40.00.0Photometric
2 February10.40.00.0Photometric
1 February7.02.41.0Photometric until close for humidity. Network interface timeouts (1.0)
 

 

January 2006

Hours spent Observing: 47.5, Weather: 247.3, Technical downtime: 37.8

Reliability and Downtime
Downtime this month has been strongly dominated by the weather.  Though it has not been as bad as last February, there has been a significant amount of snow again this year. An analysis of the technical downtime reveals that it is split roughly 50:50 between alignment of portal encoders and mundane IT related faults (hard disk failures etc). In contrast to previous months, the time lost has been in a very few full nights rather than a greater number of short interruptions. The enclosure fault in particular required an emergency site visit by staff. Design work is now ongoing to investigate if the system can be made intrinsically more robust or reoccurrence prevented simply through routine monitoring.

Engineering
There have been no major scheduled engineering works since the last report. The next is planned to take place in March and will focus primarily on the repair and recommissioning of SupIRCam.

User Support and Time Allocation
Recent decisions by the PATT and JMU Telescope Allocation Committees (TACs) have led to a significant change in telescope time allocation in the present semester (2006A).  Beginning with this semester, all proposals which have not been specifically allocated time for this and subsequent semesters have been expired.  The previous arrangement whereby once a proposal was given time, it continued to be active until either the observing programme was completed or the proposal used up all its time allocation, has been discontinued. Proposals which have been allocated time by JMU and PATT TACs are listed on the LT website allocations page.

Accordingly, please note that IF YOUR PROPOSAL IS NOT LISTED IN 2006A on the above webpages then IT HAS BEEN EXPIRED AND IS NO LONGER ACTIVE. A further application will be necessary to the appropriate TAC to re-activate the proposal in a future semester.

The decisions of the two TACs were made to make available more time on the LT and encourage competitive submission and scrutiny of existing and new proposals. Both TACs are certainly keen to receive proposals which are continuations of existing programmes.

Autoguider
We are well aware that several users have been significantly affected by the unavailability of the LT Autoguider. Over the past couple of months we have been actively investigating alternative ways of resolving the many problems we have had with providing this facility and we are pleased to be able to report we now have a way forward.

The original autoguider for the LT was supplied by TTL. This failed during the telescope commissioning process when the chip delaminated from its cooler, causing permanent damage. This month the decision has therefore been confirmed to take the process of designing and building an autoguider in-house at the ARI.  The new hardware design is based around a commercial off-the-shelf CCD camera head with a back illuminated detector. The control and centroiding software will be developed in-house at the ARI building on the demonstrated success of RATCam and RINGO instruments. As well as ensuring the work is done in a timely fashion, this also has the advantage of allowing a much tighter coupling of the autoguider control with the robotic control system of the telescope which was also developed at the ARI.

Guided by the specific requests of users, we have adopted a very aggressive schedule for the required software development and plan by June 2006 to have a working autoguider available for use on the LT. We will update you with progress as information becomes available.

Date Hours Used Comment
ObservingWeatherTechnical
31 January3.07.50.0High cloud. Non-photometric
30 January10.50.00.0Full robotic night
29 January0.00.010.5Problems with TCC disk
28 January0.010.60.0Ice
27 January0.010.60.0Humidity
26 January0.010.60.0Humidity
25 January0.010.60.0Rain
24 January0.010.60.0Rain
23 January0.06.83.8Enclosure fault
22 January0.00.010.6Enclosure fault
21 January10.60.00.0Full night robotic operations
20 January0.010.70.0Humidty, Cloud, Ice
19 January0.010.70.0Heavy Ice buildup
18 January0.010.70.0Heavy Ice buildup
17 January0.010.70.0High humidity, Ice
16 January0.010.70.0High humidity, Ice
15 January0.010.70.0High humidity
14 January0.00.010.7Altitude Node will not boot
13 January0.010.70.0High humidity
12 January0.010.80.0High humidity
11 January0.010.80.0Snow and high humidity
10 January0.010.90.0Humidity and cloud
9 January0.010.90.0Its snowing...
8 January0.010.90.0Its snowing...
7 January5.53.32.2Humidity until 11pm. Oil Temperature Fault at 04:30 am.
6 January0.010.90.0Humidity, Ice
5 January0.010.90.0Humidity, Ice
4 January0.010.90.0Freezing Fog
3 January0.010.90.0High humidity
2 January10.90.00.0Non-Photometric.
1 January7.03.90.0Photometric. High humidity spikes at start of night.